OpenStudy (anonymous):

trying to figure out how to solve this problem 2/5 + 7/10 x -3/2 + 4/15

OpenStudy (anonymous):

\[\frac{2}{5} + \frac{7}{10} * \frac{-3}{2} + \frac{4}{15}\] Find the lowest common denominator (LCD), which happens to be 30 \[\frac{12}{30} + \frac{21}{30} * \frac{-45}{30} + \frac{8}{30}\]

OpenStudy (anonymous):

Oh crap, scratch that \[\frac{2}{5} + \frac{-21}{20} + \frac{4}{15}\]

OpenStudy (anonymous):

The LCD is 60, so \[\frac{24}{60} + \frac{-63}{60} + \frac{16}{60}\]

OpenStudy (anonymous):

okay so that comes out to -55/60 right but the they say -55/60 = -11/12 I am not sure how they determined that???

OpenStudy (anonymous):

You have to multiply \[\frac{5}{5} * \frac{-11}{12}\] to get \[\frac{-55}{60}\]. Remember that \[\frac{5}{5} = 1\] and multiplying any number by 1 will equal that number.